OpenStudy (anonymous):

Will fan and medal :) maybe you can look at this shiraz? Simplify the rational expression by rationalizing the denominator.

OpenStudy (anonymous):

\[4 \sqrt{150}\div \sqrt{189x}\]

OpenStudy (anonymous):

this is what it really looks like: \[\frac{ 4 \sqrt{150} }{ \sqrt{189x} }\]

OpenStudy (shiraz14):

You need to multiply the fraction by the denominator to rationalize it. So this becomes \[\frac{ 4\sqrt{150} \sqrt{189x} }{ 189x }\] = \[\frac{ 12\sqrt{150} \sqrt{21x} }{ 189x }\] = \[\frac{ 4\sqrt{150} \sqrt{21x} }{ 63x }\] = \[\frac{ 12\sqrt{50} \sqrt{7x} }{ 63x }\] = \[\frac{ 4\sqrt{50} \sqrt{7x} }{ 21x }\] = \[\frac{ 20 \sqrt{14x} }{ 21x }\]